OpenStudy (anonymous):

Sam is observing the velocity of a car at different times. After three hours, the velocity of the car is 51 km/h. After five hours, the velocity of the car is 59 km/h. Write an equation in two variables in the standard form that can be used to describe the velocity of the car at different times. Show your work and define the variables used. How can you graph the equation obtained in Part A for the first six hours?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

you are given 2 points of reference, how do we create a line equation using 2 points?

OpenStudy (anonymous):

plot the two points and connect them

OpenStudy (amistre64):

thats how we graph them, thats not how we make an equation

OpenStudy (anonymous):

ohh, y = mx + b

OpenStudy (amistre64):

or if we dont know the y intercept: just use point and slope y = m(x-xo) + yo

OpenStudy (amistre64):

xo,yo is either point of reference, so all we need to do is determine the slope between the points

OpenStudy (anonymous):

(3, 51) and (5, 59) m = y2 - y1/x2 - x1 m = 59 - 51/5 - 3 m = 8/2 m = 4 y = mx + b y = 4x + b (3,51) y = 4x + b 51 = 4(3) + b 51 = 12 + b 51 - 12 = 39 b = 39 (5,59) y = 4x + b 59 = 4(5) + b 59 = 20 + b 59 - 20 = 39 b = 39 y = 4x + 39

OpenStudy (anonymous):

@amistre64 is this correct?

OpenStudy (amistre64):

(3, 51) and -(5, 59) -------- -2,-8: slope is is good now the equation is just: y = 4(x-3)+51 using the slope and the point (3,51)

OpenStudy (amistre64):

in your case y = 51-12 = 41-2 = 39 is finne as well

OpenStudy (amistre64):

b=, not y= had a typo

OpenStudy (amistre64):

so, this gives us a linear model for the points of reference, and we can now guess at the solution for what was it, x=6?

OpenStudy (anonymous):

y = 4(6) + 39 y = 24 + 39 y = 63

OpenStudy (amistre64):

we can know graph the equation for x=0 thru x=6,

OpenStudy (amistre64):

good: 0,39 and 6,63 are the end points for our graph

OpenStudy (anonymous):

Thank you so much for your help, I really appreciate it. I had no idea what they were asking me to do, lol.

OpenStudy (amistre64):

youre welcome